# MLLM Machine Unlearning via Visual Knowledge Distillation

### v1

@@ Abstract
Recently, machine unlearning approaches have been proposed to remove sensitive information from well-trained large models. However, most existing methods are tailored for LLMs, while MLLM-oriented unlearning remains at its early stage. Inspired by recent studies exploring the internal mechanisms of MLLMs, we propose to disentangle the visual and textual knowledge embedded within MLLMs and introduce a dedicated approach to selectively erase target visual knowledge while preserving textual knowledge. Unlike previous unlearning methods that rely on output-level supervision, our approach introduces a Visual Knowledge Distillation (VKD) scheme, which leverages intermediate visual representations within the MLLM as supervision signals. This design substantially enhances both unlearning effectiveness and model utility. Moreover, since our method only fine-tunes the visual components of the MLLM, it offers significant efficiency advantages. Extensive experiments demonstrate that our approach outperforms state-of-the-art unlearning methods in terms of both effectiveness and efficiency. Moreover, we are the first to evaluate the robustness of MLLM unlearning against relearning attacks.
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) have achieved remarkable progress, demonstrating outstanding performance on various multimodal applications such as image captioning and visual question answering (VQA) [liu2023visual, bai2023qwen]. However, growing concerns over data privacy have become a significant barrier to the widespread application of LLMs and MLLMs. This is because training large models typically involves the use of vast amounts of Internet data, which often contain sensitive personal information such as social security numbers or personal photographs. Moreover, numerous studies [pi2024mllm, li2024digger, cohen2024performance] have shown that large models can memorize portions of their training data and reproduce them verbatim in their outputs. This poses a substantial risk of privacy leakage. Therefore, the advent of data privacy and protection regulations—such as the European Union’s GDPR [regulation2018general] and the California Consumer Privacy Act (CCPA) [goldman2020introduction]—now mandates these large model providers to honor data deletion requests from individuals [dang2021right].
@@
The most straightforward approach to protecting privacy is to discard the trained model entirely, remove the individual’s data from the training set, and retrain a new model from scratch. However, retraining is computationally expensive and resource-intensive, particularly for large models. As a result, the field of machine unlearning [garg2020formalizing, ginart2019making, cao2015towards, gupta2021adaptive, sekhari2021remember, brophy2021machine, ullah2021machine] has emerged, focusing on the challenge of efficiently revising a trained model to forget specific data while preserving performance on the retaining dataset. Nowadays, many large-model-oriented machine unlearning methods have been proposed, such as Gradient Ascent (GA) [thudi2022unrollingsgd], its variations [liu2022privateunlearning], and Negative Preference Optimization (NPO) [zhang2024npo], etc. However, most existing approaches primarily focus on unlearning for LLMs, with relatively limited exploration in the context of MLLMs.
@@
In general, unlearning in LLMs is clearly defined—erasing the knowledge of a specific entity from the model. In contrast, defining unlearning for MLLMs is far more complex, as two distinct forms of knowledge are associated with the entity of interest: textual knowledge, stored within the LLM component, and visual knowledge, or visual patterns, linked to the entity’s appearance. For instance, for a well-known individual, the name (e.g., “Robin Williams”) and textual attributes (such as occupation or home address) constitute the textual knowledge, while the person’s facial appearance represents the visual knowledge. Consequently, there remains ambiguity in defining MLLM unlearning: some studies argue that it should involve erasing both textual and visual knowledge [liu2025MANU], whereas others contend that only the visual knowledge should be removed while preserving the textual knowledge [huo2025mmunlearner].
@@
In this paper, we follow [huo2025mmunlearner] in defining the desired MLLM unlearning objective as erasing visual knowledge while preserving textual knowledge, as this formulation offers greater flexibility. The essence of this definition lies in disentangling visual and textual knowledge, thereby allowing selective removal of either component when necessary. Even when the goal is to erase both textual and visual knowledge, this can be achieved by sequentially performing MLLM unlearning followed by LLM unlearning.
@@
An MLLM typically consists of three components: a vision encoder, an LLM backbone, and a projector that bridges the two. Unlike [huo2025mmunlearner], which fine-tunes all three components for unlearning, we propose to fine-tune only the vision encoder and the projector while keeping the LLM backbone frozen. This design is motivated by some studies that explore the internal mechanisms of MLLMs [cohen2024performance, huang2024commonsenseknowledgeeditingbased, yu2024understanding]. These works decompose the VQA process into two stages: the first involves visually recognizing the entity (termed identification), and the second connects the recognized entity to its stored factual knowledge (termed extraction). It has been shown that, in a well-trained MLLM, the vision encoder and projector primarily handle the identification function, whereas the LLM module is responsible for the extraction function. Therefore, to effectively erase visual knowledge while preserving textual knowledge, we propose to keep the LLM module frozen during the unlearning process.
Note that  [huo2025mmunlearner] conducted an ablation study comparing such partial fine-tuning setting (fine-tuning only the vision encoder and projector) with the full fine-tuning setting. Their results demonstrated that the former is inferior to the latter in balancing unlearning effectiveness and model utility. In this paper, we propose a novel scheme, termed Visual Knowledge Distillation (VKD), to harness the advantages of our partial fine-tuning setting.
@@
Specifically, when designing the unlearning objective function, most existing methods focus primarily on the model’s output. Take GA as an example: for the forgetting knowledge, it encourages the model to avoid producing the ground-truth answer, while for the remaining knowledge, it encourages the model to retain correct predictions. We argue that such output-level supervision is well-suited for full fine-tuning settings. However, in the partial fine-tuning scenario, this supervision signal must propagate through the frozen LLM, and thus becomes significantly weakened by the time it reaches the visual module.
@@
To address this issue, we propose to inject the supervision signal directly into the visual module. Concretely, we employ the original MLLM as the teacher model and use its intermediate visual representation (i.e., the output of the projector) as the supervision signal. We then introduce a loss term that encourages the unlearned model (the student model) to align with this signal. We term this approach Visual Knowledge Distillation (VKD), as it distills the teacher model’s visual representations—which predominantly encode visual knowledge—into the student model during unlearning. The empirical results demonstrate that our approach outperforms previous methods in terms of both unlearning effectiveness and model utility.
@@
In addition to its superior unlearning performance, our partial fine-tuning scheme also offers notable advantages in efficiency. Since the knowledge distillation is applied solely to the visual component, the fine-tuning process is highly efficient, requiring updates to only a small portion of parameters. Furthermore, due to its lightweight nature, our approach is well-suited for online unlearning scenarios, where data deletion requests occur frequently and arrive sequentially. In such cases, our method can efficiently fine-tune the visual component on demand for each request.
@@
Beyond evaluating the effectiveness of unlearning, we also examine its robustness. Recent studies have revealed that machine unlearning is vulnerable to the so-called relearning attack, which attempts to recover forgotten data or knowledge using only a few samples from the forgetting set [lynch2024robustunlearning, hu2025joggingmemory]. If the ‘already-unlearned’ knowledge can be rapidly recovered, it suggests that the unlearning process merely obfuscates the knowledge rather than truly removes it. Notably, this is the first work to consider the robustness of unlearning in MLLMs.
@@ 2 Related work
Since LLMs may unintentionally memorize their pretraining data, Machine Unlearning (MU) techniques have been introduced to erase privacy-sensitive or copyrighted information from trained models. Gradient Ascent (GA) [thudi2022unrolling] is a representative method that fine-tunes the model to reduce the probability of generating sensitive information. One inherent challenge in unlearning lies in balancing unlearning effectiveness and model utility—that is, erasing the target data while preserving the remaining data. To address this, several GA variants have been proposed to mitigate the impact on remaining data by incorporating KL-divergence minimization [maini2024tofu] or by explicitly reducing the loss on the retained data [rafailov2023direct]. Beyond supervised fine-tuning, several reinforcement learning-based methods, such as Negative Preference Optimization (NPO) [zhang2024negative], have been proposed to further enhance the effectiveness of unlearning.
@@
Compared to MU for LLMs, research on MU for MLLMs remains in its infancy. SIU [li2024single] explores the erasure of visual patterns in MLLMs using real-world entity datasets through multifaceted fine-tuning. As noted by [yu2024understanding, cohen2024performance], MLLMs exhibit strong cross-modal interactions that entangle visual and textual knowledge, making selective unlearning more challenging and often resulting in substantial unintended knowledge loss. Moreover, even the definition of MLLM unlearning remains controversial: some studies define it as erasing visual knowledge while preserving textual knowledge [huo2025mmunlearner], whereas others argue it should involve forgetting both visual and textual knowledge [liu2025modality]. Nevertheless, existing methods predominantly rely on model outputs as the supervision signal for unlearning. In contrast, our approach leverages intermediate representations through a visual knowledge distillation strategy, providing more direct and effective supervision.
Figure 1: Illustration of our machine unlearning framework for MLLMs. For the target entity, our objective is to (a) forget its visual knowledge while (b) preserving its textual knowledge. For non-target entities, we preserve both (c) the textual knowledge and (d) the visual knowledge through Visual Knowledge Distillation (VKD).
@@ 3.1 MLLM Unlearning Problem
An MLLM model $M$ typically consists of three main components: a vision encoder $\mathcal{V}$, an LLM module $\mathcal{W}$, and a projector $\mathcal{P}$ that bridges the two. Generally, an MLLM model is trained on a visual instruction dataset $D={(v_{i},q_{i},a_{i})}_{i=1}^{N}$, where $v_{i}$ denotes the input image, $q_{i}$ the textual question, and $a_{i}$ the corresponding answer. Specifically, the training of the model $M$ is carried out through a visual question answering (VQA) task as follows: (1) The image $v_{i}$ is first processed by the vision encoder and the projector, yielding the visual embedding $v^{e}_{i}=\mathcal{P}(\mathcal{V}(v_{i}))$. (2) The question $q_{i}$ is tokenized into a textual embedding $q^{e}_{i}$, which is then concatenated with $v^{e}_{i}$ and fed into the LLM module $\mathcal{W}(q^{e}_{i}\oplus v^{e}_{i})$. (3) The LLM subsequently generates an output that is expected to align with the target answer $a_{i}$.
@@ 3.2 Our Approach
Since modern MLLMs are typically built upon a pretrained LLM backbone, it is widely assumed that textual knowledge is primarily encoded in the LLM, while visual knowledge is largely captured by the visual module [zhu2023minigpt, chen2024internvl]. Moreover, one of our core unlearning goals is to preserve textual knowledge while removing visual knowledge. Therefore, freezing the parameters of the LLM naturally benefits the retention of textual knowledge.
@@
3.2.1 Visual Knowledge Distillation
@@
Notably, the term $f^{r}_{vqa}-t^{r}_{vqa}||^{2}$ encourages the unlearned model to preserve visual knowledge unrelated to the target entity. Our empirical results indicate that when only the visual module is fine-tuned, the key to balancing unlearning effectiveness and model utility lies in retaining the visual knowledge unrelated to the target entity, while the erasure of target visual knowledge can be sufficiently achieved through output-level supervision.
Our approach involves two weighting parameters. Specifically, $\beta$ is introduced to balance the output-level objective and the VKD objective. The parameter $\alpha$ is employed to trade off among the visual knowledge to be forgotten ($\mathcal{L}^{f}_{vqa}$), and the visual and textual knowledge to be preserved ($\mathcal{L}^{r}_{vqa}$, $\mathcal{L}^{f}_{qa}$ and $\mathcal{L}^{r}_{qa}$).
3.2.2 Selective Forgetting
@@
As shown in Eq 4, we follow the method [huo2025mmunlearner] that leverage a masking scheme as performing unlearning, where m is a boolean mask that selectively updates parameters, and $\odot$ denotes the Hadamard product. Specifically, the mask m is referred to as the weight saliency map, as it quantifies how salient each parameter is with respect to a given dataset $D$.
In addition to saliency-based masking, another selective forgetting technique is neuron pruning, which identifies neurons closely associated with the knowledge to be forgotten and achieves unlearning by pruning them. In this paper, we integrate these two selective forgetting strategies. They can be executed either in parallel or sequentially. Our experiments demonstrate that the best performance is achieved when pruning is applied first, followed by masking-based fine-tuning.
@@
It is worth noting that in our approach, both neuron pruning and weight masking are applied exclusively to the visual module—specifically, the vision encoder $\mathcal{V}$ and the projector $\mathcal{P}$. This is motivated by recent studies on localizing factual memory within Transformer architectures. Prior research [liu2025MANU] has shown that MLP layers serve as primary knowledge storage components in Transformer architectures [geva2021transformerfeedforwardlayerskeyvalue, meng2023locatingeditingfactualassociations], whereas the self-attention modules mainly function as mechanisms for information propagation [elhage2021mathematical].
@@
In particular, for vision encoders such as the CLIP model, it has been observed that the high-dimensional inner projections within the feed-forward layers play a crucial role in storing visual knowledge, whereas the representations derived from the self-attention layers do not [ghiasi2022visiontransformerslearnvisual]. Furthermore, it has been also observed that earlier layers learn textural attributes, whereas deeper layers learn high level object features or abstract concepts [ghiasi2022visiontransformerslearnvisual]. Therefore, in our approach, both neuron pruning and weight masking are applied exclusively to the vision encoder’s MLP layers, with particular emphasis on the deeper layers.
4 Evaluation
4.1 Experimental Setup
To accurately assess whether the target knowledge has been successfully unlearned from the MLLM, we typically do not perform unlearning directly on the original pretrained model (e.g., LLaVA-1.5-7B), as it is uncertain whether the model initially possesses the target knowledge. In practice, most unlearning benchmarks adopt a two-stage procedure: first, a fictitious dataset representing the target knowledge (e.g.,, synthetic personal profiles) is curated, and the original MLLM is fine-tuned on this dataset to obtain a vanilla model that explicitly contains the target knowledge. Then, the unlearning algorithm is applied to this vanilla model, and the resulting unlearned model is evaluated to assess the effectiveness of the unlearning process.
@@ 4.1 Experimental Setup
MMU 44.0 54.4 56.0 55.7 75.3 77.3 50.0 0.123 70.9 0.100 68.9 94.7
Ours 42.8 54.6 56.0 56.4 74.9 77.4 48.7 0.125 71.5 0.094 69.0 90.8
Table 1: Overall results of baselines and our approach on MLLMU-Bench and CLEAR benchmarks. $\downarrow$ Lower is better for Forget VQA accuracy (reflecting better forgetting of target visual knowledge), $\uparrow$ Higher is better for the others (reflecting better retention of non-target visual knowledge and textual knowledge). For MLLMU-Bench, the Forget/Retain QA tasks are evaluated using the Acc metric, while for CLEAR, these tasks are evaluated using the ROUGE-L metric.
4.1.1 Datasets and Metrics
Datasets: We employ two benchmark datasets to evaluate unlearning performance in MLLMs. MLLMU-Bench comprises fictitious personal profiles, each paired with a portrait and 18 corresponding questions (i.e., 9 VQA questions and 9 textual QA questions) with multiple-choice options. The dataset is divided into three subsets: Forget, Retain, and Real-world. The Forget Set contains a subset of fictitious profiles designated for evaluating unlearning effectiveness. The Retain Set includes fictitious profiles excluded from the Forget Set, representing the knowledge the model should preserve. The Real-world Set consists of real celebrity profiles distinct from the fictitious ones and is used to assess unlearning entanglement with neighboring real-world concepts, thereby reflecting the model’s utility. CLEAR is built on top of an LLM unlearning dataset, i.e., the TOFU dataset. For each author in TOFU, CLEAR adds several face images and corresponding textual descriptions generated by GPT-4o [achiam2023gpt]. Similar to MLLMU-Bench, CLEAR also comprises Forget, Retain, and Real-world sets.
Metrics: We evaluate the forgetting and retention of textual knowledge using the Question Answering (QA) task, and assess the forgetting and retention of visual knowledge using the Visual Question Answering (VQA) task. The performance on multiple-choice questions is measured by average accuracy, while text generation performance is evaluated using the ROUGE-L metric [lin2004rouge].

### v2

# Robust MLLM Unlearning via Visual Knowledge Distillation

@@ Abstract
Recently, LLM unlearning approaches have been proposed to remove sensitive information from well-trained large models. However, unlearning for Multimodal Large Language Models (MLLMs) remains at an early stage. Inspired by recent studies on the internal mechanisms of MLLMs, we propose to disentangle visual and textual knowledge within MLLMs and introduce a dedicated approach that selectively erases target visual knowledge while preserving textual knowledge. Unlike previous unlearning methods that rely on output-level supervision, our approach introduces a Visual Knowledge Distillation (VKD) scheme, which leverages intermediate visual representations within the MLLM as supervision signals. This design substantially enhances both unlearning effectiveness and model utility. Moreover, since our method only fine-tunes the visual components of MLLM, it offers significant efficiency advantages. Extensive experiments demonstrate that our approach outperforms state-of-the-art unlearning methods in terms of both effectiveness and efficiency. Furthermore, we are the first to evaluate the robustness of MLLM unlearning against re-learning attacks.
Machine Learning, ICML
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) have achieved remarkable progress. However, growing concerns over data privacy have become a significant barrier to the widespread application of LLMs and MLLMs. This is because training large models typically involves the use of vast amounts of Internet data, which often contain sensitive personal information such as social security numbers or personal photographs. Moreover, numerous studies (Pi et al., 2024; Li et al., 2024a; Cohen et al., 2024) have shown that large models can memorize portions of their training data and reproduce them verbatim in their outputs. This poses a substantial risk of privacy leakage. Therefore, the advent of data privacy and protection regulations—such as the European Union’s GDPR (Regulation, General Data Protection, 2018) and the California Consumer Privacy Act (CCPA) (Goldman, 2020)—now mandates these large model providers to honor data deletion requests from individuals (Dang, 2021).
@@
The most straightforward approach to protecting privacy is to discard the trained model entirely, remove the individual’s data from the training set, and retrain a new model from scratch. However, retraining is computationally expensive and resource-intensive, particularly for large models. As a result, the field of machine unlearning (Garg et al., 2020; Ginart et al., 2019; Cao and Yang, 2015; Gupta et al., 2021; Sekhari et al., 2021; Brophy and Lowd, 2021; Ullah et al., 2021) has emerged, aiming to efficiently revise trained models by selectively forgetting specific data while preserving performance on the remaining data. Nowadays, many LLM unlearning methods have been proposed, such as Gradient Ascent (GA) (Thudi et al., 2022), its variations (Liu et al., 2022), and Negative Preference Optimization (NPO) (Zhang et al., 2024), etc. However, MLLM unlearning remains at an early stage.
@@
LLM unlearning problem is clearly defined—erasing the knowledge of a specific entity from the model. In contrast, defining unlearning for MLLMs is far more complex, as two distinct forms of knowledge are associated with the entity of interest: textual knowledge, linked to the entity’s general information, and visual knowledge/visual patterns, linked to the entity’s appearance. For instance, for a well-known individual, the name (e.g., “Robin Williams”) and textual attributes (such as occupation or home address) constitute the textual knowledge, while the person’s facial appearance represents the visual knowledge. Consequently, there remains ambiguity in defining MLLM unlearning: some studies argue that it should involve erasing both textual and visual knowledge (Liu et al., 2025), whereas others contend that only the visual knowledge should be removed while preserving the textual knowledge (Huo et al., 2025).
@@
In this paper, we follow (Huo et al., 2025) to define the MLLM unlearning objective as erasing visual knowledge while preserving textual knowledge, as this formulation offers greater flexibility. The essence of this definition lies in disentangling visual and textual knowledge, thereby allowing selective removal of either component when necessary. Even when the goal is to erase both textual and visual knowledge, this can be achieved by sequentially performing MLLM unlearning followed by LLM unlearning.
@@
An MLLM typically consists of three components: a vision encoder, an LLM backbone, and a projector that bridges the two. Unlike (Huo et al., 2025), which fine-tunes all three components for unlearning, we propose to fine-tune only the vision encoder and the projector while keeping the LLM backbone frozen. This design is motivated by some studies that explore the internal mechanisms of MLLMs (Cohen et al., 2024; Huang et al., 2024a; Yu and Ananiadou, 2024). These works decompose the VQA process into two stages: the first involves visually recognizing the entity (termed identification), and the second connects the recognized entity to its stored factual knowledge (termed extraction). It has been shown that, in a well-trained MLLM, the vision encoder and projector primarily handle the identification function, whereas the LLM module is responsible for the extraction function. Therefore, fine-tuning the visual module and disrupting the identification stage is sufficient to erase visual knowledge. Moreover, freezing the LLM component can largely preserve textual knowledge.
@@
Note that  (Huo et al., 2025) conducted an ablation study comparing such partial fine-tuning scheme (fine-tuning only the vision encoder and projector) with the full fine-tuning scheme. Their results demonstrated that the former is inferior to the latter in balancing unlearning effectiveness and model utility. In this paper, we propose a novel approach, termed Visual Knowledge Distillation (VKD), to harness the advantages of partial fine-tuning scheme. Specifically, when designing the unlearning objective function, most existing methods focus primarily on the model’s output. Take GA as an example: for the forgetting knowledge, it encourages the model to avoid producing the ground-truth answer, while for the remaining knowledge, it encourages the model to retain correct predictions. We argue that such output-level supervision is well-suited for full fine-tuning scheme. However, in the partial fine-tuning scenario, this supervision signal must propagate through the frozen LLM, and thus becomes significantly weakened by the time it reaches the visual module.
@@
To address this issue, our VKD approach injects the supervision signal directly into the visual module. Concretely, we employ the original MLLM as the teacher model and use its intermediate visual representation (i.e., the output of the projector) as the supervision signal. We then introduce a loss term that encourages the unlearned model (the student model) to align with this signal. Our VKD distills the teacher model’s intermediate visual representations—which predominantly encode visual knowledge—into the student model during unlearning. The empirical results demonstrate that our approach outperforms previous methods in terms of both unlearning effectiveness and model utility.
@@
In addition to its superior unlearning performance, our partial fine-tuning scheme also offers notable advantages in efficiency. Since the knowledge distillation is applied solely to the visual component, the fine-tuning process is highly efficient, requiring updates to only a small portion of parameters. Furthermore, due to its lightweight nature, our approach is well-suited for continual unlearning scenarios, where data deletion requests occur frequently and arrive sequentially. In such cases, our method can efficiently fine-tune the visual component on demand for each request.
@@
Beyond evaluating the effectiveness of unlearning, we also examine its robustness. Recent studies have revealed that machine unlearning is vulnerable to the so-called relearning attack, which attempts to recover forgotten data or knowledge using only a few samples from the forgetting set (Lynch et al., 2024; Hu et al., 2025). If the ‘already-unlearned’ knowledge can be rapidly recovered, it suggests that the unlearning process merely obfuscates the knowledge rather than truly removes it. Notably, this is the first work to consider the robustness of unlearning in MLLMs.
@@ 2 Related Work
Since LLMs may unintentionally memorize their pretraining data, Machine Unlearning (MU) techniques have been introduced to erase privacy-sensitive or copyrighted information from trained models. Gradient Ascent (GA) (Thudi et al., 2022) is a representative method that fine-tunes the model to reduce the probability of generating sensitive information. One inherent challenge in unlearning lies in balancing unlearning effectiveness and model utility—that is, erasing the target data while preserving the remaining data. To address this, several GA variants have been proposed to mitigate the impact on remaining data by incorporating KL-divergence minimization (Maini et al., 2024) or by explicitly reducing the loss on the retained data (Rafailov et al., 2023). Beyond supervised fine-tuning, several reinforcement learning-based methods, such as Negative Preference Optimization (NPO) (Zhang et al., 2024), have been proposed to further enhance the effectiveness of unlearning.
@@
Compared to MU for LLMs, research on MU for MLLMs remains in its infancy. SIU (Li et al., 2024b) explores the erasure of visual patterns in MLLMs using real-world entity datasets through multifaceted fine-tuning. As noted by (Yu and Ananiadou, 2024; Cohen et al., 2024), MLLMs exhibit strong cross-modal interactions that entangle visual and textual knowledge, making selective unlearning more challenging and often resulting in substantial unintended knowledge loss. Moreover, even the definition of MLLM unlearning remains controversial: some studies define it as erasing visual knowledge while preserving textual knowledge (Huo et al., 2025), whereas others argue it should involve forgetting both visual and textual knowledge (Liu et al., 2025). Nevertheless, existing methods predominantly rely on model outputs as the supervision signal for unlearning. In contrast, our approach leverages intermediate representations through a visual knowledge distillation strategy, providing more direct and effective supervision.
Figure 1: Illustration of our machine unlearning framework for MLLMs. For the target entity, our objective is to (a) forget its visual knowledge while (b) preserving its textual knowledge. For non-target entities, we preserve both (c) the textual knowledge and (d) the visual knowledge through VKD.
@@ 3.1 MLLM Unlearning Problem
An MLLM model $M$ typically consists of three main components: a vision encoder $\mathcal{V}$, an LLM module $\mathcal{W}$, and a projector $\mathcal{P}$ that bridges the two. Generally, an MLLM model is trained on a visual instruction dataset $D={(v_{i},q_{i},a_{i})}_{i=1}^{N}$, where $v_{i}$ denotes the input image, $q_{i}$ the textual question, and $a_{i}$ the corresponding answer.
Specifically, the training of the model $M$ is carried out through a visual question answering (VQA) task as follows: (1) The image $v_{i}$ is first processed by the vision encoder and the projector, yielding the visual embedding $v^{e}_{i}=\mathcal{P}(\mathcal{V}(v_{i}))$. (2) The question $q_{i}$ is tokenized into a textual embedding $q^{e}_{i}$, which is then concatenated with $v^{e}_{i}$ and fed into the LLM module $\mathcal{W}(q^{e}_{i}\oplus v^{e}_{i})$. (3) The LLM subsequently generates an output that is expected to align with the target answer $a_{i}$.
@@ 3.2 Our Approach
Since modern MLLMs are typically built upon a pretrained LLM backbone, it is widely assumed that textual knowledge is primarily encoded in the LLM, while visual knowledge is largely captured by the visual module (Zhu et al., 2023; Chen et al., 2024). Moreover, one of our core unlearning goals is to preserve textual knowledge while removing visual knowledge. Therefore, freezing the parameters of the LLM naturally benefits the retention of textual knowledge.
@@
Visual Knowledge Distillation.
@@
Notably, the term $f^{r}_{vqa}-t^{r}_{vqa}||^{2}$ encourages the unlearned model to preserve visual knowledge unrelated to the target entity. Our empirical results indicate that when only the visual module is fine-tuned, the key to balancing unlearning effectiveness and model utility lies in retaining the visual knowledge unrelated to the target entity, while the erasure of target visual knowledge can be sufficiently achieved through output-level supervision. Our approach involves two weighting parameters. Specifically, $\beta$ is introduced to balance the output-level objective and the VKD objective. The parameter $\alpha$ is employed to trade off among the visual knowledge to be forgotten ($\mathcal{L}^{f}_{vqa}$), and the visual and textual knowledge to be preserved ($\mathcal{L}^{r}_{vqa}$, $\mathcal{L}^{f}_{qa}$ and $\mathcal{L}^{r}_{qa}$).
Another advantage of intermediate-level supervision is that it eliminates the need to load the full reference model into memory; instead, only the visual module is required to obtain the supervision signal.
Selective Forgetting.
@@
As shown in Eq 5, we follow the method (Huo et al., 2025) that leverage a masking scheme as performing unlearning, where m is a boolean mask that selectively updates parameters, and $\odot$ denotes the Hadamard product. Specifically, the mask m is referred to as the weight saliency map, as it quantifies how salient each parameter is with respect to a given dataset $D$. In addition to saliency-based masking, another selective forgetting technique is neuron pruning (Pochinkov and Schoots, 2024), which identifies neurons associated with the knowledge to be forgotten and achieves unlearning by pruning them.
In this paper, we integrate these two selective forgetting strategies into a unified framework. There are multiple ways to combine them, such as parallel or sequential integration. Our experiments demonstrate that the best performance is achieved when pruning is applied first, followed by masking-based fine-tuning. Specifically, we adopt the neuron pruning strategy based on importance functions, as proposed in (Pochinkov and Schoots, 2024). We leverage the mean of absolute activation $I_{abs}$ as the importance function. Let $n$ be a neuron and denote its corresponding activations by $z$, the importance of neuron $n$ with respect to the dataset $D$ is defined as follows:
$I_{\text{abs}}(D,n)=\frac{1}{|D|}\sum_{i\in D}|z(i)|$ (6)
where $i\in D$ indicates the data sample $i$ in $D$.
Thus, the neuron importance for unlearning $I_{u}$ is derived by the ratio of $I_{abs}$ between the knowledge to be forgotten and the knowledge to be retained:
$I_{u}=\frac{I_{abs}(D^{f}_{vqa})}{I_{abs}(D^{f}_{qa})+I_{abs}(D^{r}_{vqa})+I_{abs}(D^{r}_{qa})}$ (7)
Consequently, we identify the neurons to be pruned, denoted as $\textbf{n}_{P}$, by applying a threshold to $I_{u}$:
$\textbf{n}_{P}=\mathbb{I}[I_{u}\geq d_{I}]$ (8)
where $d_{I}$ is a predefined threshold.
After selective neuron pruning, we perform masking-based fine-tuning. The key lies in computing the saliency of each model weight with respect to the knowledge to be forgotten. Previous studies (Huang et al., 2024b) have shown that this saliency can be approximated using the diagonal elements of the Fisher Information Matrix (FIM) of the vanilla model with respect to dataset $D$:
$F^{D}_{diag}=[\Delta\mathcal{L}^{D}(\theta_{0})]^{2}$ (9)
where $\theta_{0}$ denotes the vanilla model.
Thus, the following ratio R indicates which parameters are salient for forgetting the target visual knowledge while preserving other knowledge:
$R=\frac{[\Delta\mathcal{L}^{f}_{vqa}]^{2}}{[\Delta(\mathcal{L}^{f}_{qa}+\mathcal{L}^{r}_{vqa}+\mathcal{L}^{r}_{qa})]^{2}}$ (10)
The Fisher mask $\textbf{m}_{F}$ is then obtained by comparing the ratio R against a predefined threshold $d_{F}$:
$\textbf{m}_{F}=\mathbb{I}[R\geq d_{F}]$ (11)
This mask is regarded as the final mask m used in Eq. 5.
Besides this pruning-then-fine-tuning scheme, other variant strategies are discussed in Appendix.
Unlearning on MLP layers.
@@
It is worth noting that in our approach, both neuron pruning and weight masking are applied to the visual module. This is motivated by recent studies on localizing factual memory within Transformer architectures. Prior research (Liu et al., 2025) has shown that MLP layers serve as primary knowledge storage components in Transformer architectures (Geva et al., 2021; Meng et al., 2023), whereas the self-attention modules mainly function as mechanisms for information propagation (Elhage et al., 2021).
@@
In particular, for vision encoders such as the CLIP model, it has been observed that the high-dimensional inner projections within the feed-forward layers play a crucial role in storing visual knowledge, whereas the representations derived from the self-attention layers do not (Ghiasi et al., 2022). Furthermore, it has been observed that earlier layers learn textural attributes, whereas deeper layers learn high level object features or abstract concepts (Ghiasi et al., 2022). Therefore, in our approach, both neuron pruning and weight masking are applied to the vision encoder’s MLP layers, with emphasis on the deeper layers.
@@
MMU 44.0 54.4 52.6 55.7 75.3 77.3 50.0 0.123 68.3 0.100 68.9 94.7
Ours 42.8 54.6 56.0 56.4 75.4 77.4 48.7 0.125 70.9 0.094 69.0 90.8
Table 1: Overall results of baselines and our approach on MLLMU-Bench and CLEAR benchmarks. Bold and underlined values indicate the best and second-best performance, respectively. $\downarrow$ Lower is better for Forget VQA accuracy, $\uparrow$ Higher is better for the others. For MLLMU-Bench, the Forget/Retain QA tasks are evaluated using the Acc metric, while for CLEAR, these tasks are evaluated using the ROUGE-L metric.
4 Evaluation
To accurately assess whether the target knowledge has been successfully unlearned from the MLLM, we typically do not perform unlearning directly on the original pretrained model (e.g., LLaVA-1.5-7B), as it is uncertain whether the model initially possesses the target knowledge. In practice, most unlearning benchmarks adopt a two-stage procedure: first, a fictitious dataset representing the target knowledge (e.g., synthetic personal profiles) is curated, and the base MLLM is fine-tuned on this dataset to obtain a vanilla model that explicitly contains the target knowledge. Then, the unlearning algorithm is applied to this vanilla model, and the resulting unlearned model is evaluated to assess the effectiveness of the unlearning algorithm.
4.1 Datasets and Metrics
Datasets: We employ two benchmark datasets to evaluate unlearning performance in MLLMs. MLLMU-Bench comprises fictitious personal profiles, each paired with a portrait and 18 corresponding questions (i.e., 9 VQA questions and 9 textual QA questions) with multiple-choice options. The dataset is divided into three subsets: Forget, Retain, and Real-world. The Forget Set contains a subset of fictitious profiles designated for evaluating unlearning effectiveness. The Retain Set includes fictitious profiles excluded from the Forget Set, representing the knowledge the model should preserve. The Real-world Set consists of real celebrity profiles distinct from the fictitious ones and is used to assess unlearning entanglement with neighboring real-world concepts, thereby reflecting the model’s utility. CLEAR is built on top of an LLM unlearning dataset, i.e., the TOFU dataset. For each author in TOFU, CLEAR adds several face images and corresponding textual descriptions generated by GPT-4o (Achiam et al., 2023). Similar to MLLMU-Bench, CLEAR also comprises Forget, Retain, and Real-world sets.
Metrics: We evaluate the forgetting and retention of textual knowledge using the Question Answering (QA) task, and assess the forgetting and retention of visual knowledge using the Visual Question Answering (VQA) task. The performance on multiple-choice questions is measured by average accuracy, while text generation performance is evaluated using the ROUGE-L metric (Lin, 2004).